# Instruct-ReID: A Multi-purpose Person Re-identification Task with Instructions

### v4

@@ Abstract
Human intelligence can retrieve any person according to both visual and language descriptions. However, the current computer vision community studies specific person re-identification (ReID) tasks in different scenarios separately, which limits the applications in the real world. This paper strives to resolve this problem by proposing a new instruct-ReID task that requires the model to retrieve images according to the given image or language instructions. Our instruct-ReID is a more general ReID setting, where existing 6 ReID tasks can be viewed as special cases by designing different instructions. We propose a large-scale OmniReID benchmark and an adaptive triplet loss as a baseline method to facilitate research in this new setting. Experimental results show that the proposed multi-purpose ReID model, trained on our OmniReID benchmark without fine-tuning, can improve +0.5%, +0.6%, +7.7% mAP on Market1501, MSMT17, CUHK03 for traditional ReID, +6.4%, +7.1%, +11.2% mAP on PRCC, VC-Clothes, LTCC for clothes-changing ReID, +11.7% mAP on COCAS+ real2 for clothes template based clothes-changing ReID when using only RGB images, +24.9% mAP on COCAS+ real2 for our newly defined language-instructed ReID, +4.3% on LLCM for visible-infrared ReID, +2.6% on CUHK-PEDES for text-to-image ReID. The datasets, the model, and code will be available at https://github.com/hwz-zju/Instruct-ReID.
@@ 1 Introduction
Identifying individuals exhibiting significant appearance variations from multimodal descriptions is a fundamental aspect of human intelligence with broad applications [34, 31]. To imbue our machine with this capability, person re-identification (ReID) [67, 66] has been introduced to retrieve images of the target person from a vast repository of surveillance videos or images across locations and time [10]. In recent years, significant advancements have been made in developing precise and efficient ReID algorithms and establishing benchmarks covering various scenarios, such as traditional ReID [65, 6, 47], clothes-changing ReID (CC-ReID) [18, 12, 20, 41, 15], clothes template based clothes-changing ReID (CTCC-ReID) [57, 28], visible-infrared ReID (VI-ReID) [32, 50, 51, 59] and text-to-image ReID (T2I-ReID) [3, 4, 68, 26, 56]. However, focusing solely on one specific scenario, specialized ReID models possess inherent limitations. For instance, customers must deploy distinct models to retrieve persons according to the query information, which significantly increases the cost for model training and results in inconvenience for the application. To facilitate real-world deployment, there is a pressing need to devise one generic framework capable of re-identifying individuals across all scenarios mentioned above.
@@
This paper proposes a new multi-purpose instruct-ReID task where existing 6 ReID settings can be formulated as its special cases. Specifically, the instruct-ReID uses query images and multimodal instructions as the model inputs, requiring the model to retrieve the same identity images from the gallery following the instructions. Using language or image instructions, models trained on the instruct-ReID task can be specialized to tackle diverse ReID tasks (Fig. 1a). For example, the clothes-changing ReID can be viewed as using the instruction “Ignore clothes” to retrieve. As another example, clothes template-based clothes-changing ReID can utilize a clothes template image as instruction. The proposed instruct-ReID task offers three significant advantages: easy deployment, improved performance, and easy extension to new ReID tasks. First, it enables cost-effective and convenient deployment in real-world applications. Unlike existing ReID approaches limited to specific tasks, instruct-ReID allows for utilizing a single model for various ReID scenarios, which is more practical in real applications. Second, as all existing ReID tasks can be considered special cases of instruct-ReID, we can unify the training sets of these tasks to exploit the benefits of more data and diverse annotations across various tasks – leading to enhanced performance. Third, instruct-ReID addresses a new language-instructed ReID task which requires the model to retrieve persons following language instructions. This setting is very practical in real applications for it enables customers to retrieve images they are particularly interested in through language descriptions. For example, customers can retrieve a woman wearing a black coat using one of her pictures and language instructions “Wear black coat”.
@@
To facilitate research in instruct-ReID, we introduce a new benchmark called OmniReID derived from 12 datasets  [^1] The discussion on ethical risks is provided in supplementary materials. of 6 different ReID tasks. The OmniReID benchmark exhibits two appealing characteristics. First, it emphasizes diversity by incorporating images from various domains, including surveillance scenarios and movie clips. The diversity ensures that the trained models are robust and can effectively handle ReID tasks in various real-world scenarios. Second, the OmniReID benchmark achieves comprehensiveness by offering evaluation datasets to assess various ReID tasks comprehensively, which facilitates evaluating the generalization ability of diverse ReID methods.
@@
Based on OmniReID, we design a method with the proposed adaptive triplet loss for instruct-ReID. The typical triplet loss [39] only defines positive/negative pairs by identities, failing to align with instructions. Therefore, we propose a novel adaptive triplet loss to learn a metric space that preserves identity and instruction similarities. Specifically, we design an adaptive margin between two query-instruction pairs based on instruction similarities to pull features with similar instructions close and push features with different instructions apart. This loss incorporates the instruction information into the features representation and optimizes the model to learn a metric space where similar instructions lead to closer features of query-instance pairs.
@@
In summary, the contributions of this paper are three folds. (1) We propose a new instruct-ReID task, where existing traditional ReID, clothes-changing ReID, clothes template based clothes-changing ReID, visible-infrared ReID, text-to-image ReID and language-instructed ReID can be viewed as special cases. (2) To facilitate research on instruct-ReID, we establish a large-scale and comprehensive OmniReID benchmark consisting 12 publicly available datasets. (3) We propose an adaptive triplet loss to supervise the feature distance of two query-instruction pairs to consider identity and instruction alignments. Our method consistently improves previous models on 10 datasets of 6 ReID tasks using ImageNet pretraining or ReID-specific pretraining [69]. For example, when using ReID-specific pretraining, our method improves +7.7%, +0.6%, +0.5% mAP on CUHK03, MSMT17, Market1501 for traditional ReID, +6.4%, +11.2%, +7.1% mAP on PRCC, LTCC, VC-Clothes for clothes-changing ReID when using RGB images only, +11.7% mAP on COCAS+ real2 for clothes template based clothes changing ReID, +4.3% mAP on LLCM for visible-infrared ReID, +2.6% mAP on CUHK-PEDES for text-to-image ReID, +24.9% mAP on COCAS+ real2 for the new language-instructed ReID.
@@ 2 Related work
Person Re-identification. Person re-identification aims to retrieve the same images of the same identity with the given query from the gallery set. To support the ReID task on all-weather application, various tasks are conducted on the scenarios with changing environments, perspectives, and poses [55, 6, 66, 4, 68]. Traditional ReID mainly focuses on dealing with indoor/outdoor problems when the target person wears the same clothes. Recently, to extend the application scenarios, clothes-changing ReID (CC-ReID) [57] and clothes template based clothes-changing ReID (CTCC-ReID) [28] are proposed. While CC-ReID forces the model to learn clothes-invariant features [18, 12], CCTC-ReID further extracts clothes template features [28] to retrieve the image of the person wearing template clothes. To capture person’s information under low-light environments, visible-infrared person ReID (VI-ReID) methods [59, 50, 51, 59] retrieve the visible (infrared) images according to the corresponding infrared (visible) images. In the absence of the query image, GNA-RNN [26] introduced the text-to-image ReID (T2I-ReID) task, which aims at retrieving the person from the textual description. However, existing tasks and methods focus on a single scenario, making it difficult to address the demands of cross-scenario tasks. In this paper, we introduce a new Instruct-ReID task, which can be viewed as a superset of the existing ReID tasks by incorporating instruction information into identification.
@@
Instruction Tuning. Instruction Tuning was first proposed in natural language processing (NLP) to enable large language models (LLM) to execute specific tasks by following natural language instructions. LLM instruction-tuned models, e.g., FLAN-T5 [7], InstuctGPT [36]/ChatGPT [35], can effectively prompt the ability on zero- and few-shot transfer tasks. A few works have borrowed the idea from language to vision. Flamingo [2], BLIP-2 [23], and KOSMOS-1 [17] learning with image-text pairs also show promising generalization on visual understanding tasks. While these LLM-based methods aim to generate convincing language responses following the image or language instructions, we focus on retrieving the correct person following the given instructions by tuning a vision transformer.
@@
Multimodal Retrieval. Multimodel retrieval is a widely used approach that aligns information from multiple modalities to improve model performance of the model and broaden the scope of applications. In multimodel retrieval, unimodal encoders always encode different modalities for retrieval tasks. For instance, CLIP [38], VideoCLIP [49], COOT [11] and MMV [1] utilize contrastive learning for pre-training. Other techniques like HERO [24], Clipbert [21], Vlm [48], bridgeformer [9] and UniVL [33] focus on merging different modalities for retrieval tasks to learn a generic representation. Although there have been numerous studies on multimodal retrieval, most are concentrated on language-vision pretraining or video retrieval, leaving the potential of multimodal retrieval for person ReID largely unexplored. This paper aims to investigate this underexplored area to retrieve anyone with information extracted from multiple modalities.
@@ 3 OmniReID Benchmark
Language Annotation Generation. To generate language instruction for language-instructed ReID, we annotate COCAS+ Real1 and COCAS+ Real2 with language description labels. Similar to Text-to-Image ReID datasets, e.g., CUHK-PEDES [26], language annotations in OmniReID are several sentences that describe the visual appearance of pedestrians. We divide our annotation process into pedestrian attribute generation and attribution-to-language transformation.
@@ 4 IRM: Instruct ReID Model
The proposed Instruct ReID model is shown in Fig. 3 (a). Our model consists of four parts: instruction generation (Sec. 4.1) for different ReID tasks, an editing transformer $\mathcal{E}_{e}$ (Sec. 4.2) and an instruction encoder $\mathcal{E}_{i}$ which is a visual language model, and a cross-model attention module $\mathcal{E}_{f}$. Given an instruction $\mathbf{T}$ associated with a query image $\mathbf{I}$, we obtain instruction features $\mathbf{F}_{\mathbf{T}}$ using the instruction encoder $\mathcal{E}i$. These extracted instruction features $\mathbf{F}_{T}$, along with query image tokens, are fed into the designed editing transformer $\mathcal{E}_{e}$ to obtain features $\mathbf{F}$ edited by instructions. Furthermore, we introduce an attention module $\mathcal{E}_{f}$ to efficiently combine features of the query image and instruction through the attention mechanism in the stacked transformer layers. Finally, the overall loss function is introduced (Sec. 4.4) for all ReID tasks, which includes our newly proposed adaptive triplet loss $\mathcal{L}_{atri}$ (Sec. 4.3) instead of traditional triplet loss.
@@ 4.1 Instruction Generation
Traditional ReID: Following Instruct-BLIP [8], we generate 20 instructions [^3] See all available instruction sentences in supplementary materials. from GPT-4 and randomly select one, e.g., “Do not change clothes.”, when training. The model is expected to retrieve images of the same person without altering image attributes, such as clothing. ### Query image:{Query image}
### Instruction: “Do not change clothes.”${}^{3}$
@@
[…]
### Instruction: “With clothes changed.”${}^{3}$
### Target image:{Output}
Clothes template based clothes-changing ReID: The instruction is a clothes template for a query image while a cropped clothes image for a target image. The model should retrieve images of the same person wearing the provided clothes. We provide more examples for training and test in the supplementary materials. ### Query image:{Query image}
### Instruction:{Any clothes template image}
### Target image:{Output}
[…]
### Instruction: “Retrieve cross modality image.”${}^{3}$
### Target image:{Output}
Text-to-Image ReID: The instruction is the describing sentences, and both images and text are fed into IRM during the training process. While in the inference stage, the image features and instruction features are extracted separately.  [^4] Image and instruction features are extracted separately in test stage. ### Image:{Image}${}^{4}$
### Instruction:{Sentences describing pedestrians}${}^{4}$
### Target image:{Output}
Language-instructed ReID: The instruction is several sentences describing pedestrian attributes. We randomly select the description languages from the person images in gallery and provide to query images as instruction. The model is required to retrieve images of the same person described in the provided sentences. We provide more examples for training and test in the supplementary materials. ### Query image:{Query image}
@@ 4.2 Editing Transformer
where $\mathbf{F}_{\mathbf{T}}=\mathcal{E}_{i}(\mathbf{I})$ is the instruction feature extracted by the instruction encoder $\mathcal{E}_{i}$ and $\mathbf{F}_{l-1}$ is the output feature of $(l\!-\!1)$-th zero-init transformer layer. Here, we define the initial input $\mathcal{F}_{0}$ of the first layer as $\mathbf{F}_{0}=[\mathbf{f}_{0}^{\text{CLS}},\mathbf{f}_{0}^{1},\mathbf{f}_{0}^{2},...,\mathbf{f}_{0}^{N}],$ where $\mathbf{f}^{\text{CLS}}_{0}$ is the appended [CLS] token, $(\mathbf{f}_{0}^{1},\mathbf{f}_{0}^{2},...,\mathbf{f}_{0}^{N})$ are the patch tokens of the query image and $N$ is the number of patches of the query image.

### v5

@@ Abstract
Human intelligence can retrieve any person according to both visual and language descriptions. However, the current computer vision community studies specific person re-identification (ReID) tasks in different scenarios separately, which limits the applications in the real world. This paper strives to resolve this problem by proposing a new instruct-ReID task that requires the model to retrieve images according to the given image or language instructions. Our instruct-ReID is a more general ReID setting, where existing 6 ReID tasks can be viewed as special cases by designing different instructions. We propose a large-scale OmniReID benchmark and an adaptive triplet loss as a baseline method to facilitate research in this new setting. Experimental results show that the proposed multi-purpose ReID model, trained on our OmniReID benchmark without fine-tuning, can improve +0.5%, +0.6%, +7.7% mAP on Market1501, MSMT17, CUHK03 for traditional ReID, +6.4%, +7.1%, +11.2% mAP on PRCC, VC-Clothes, LTCC for clothes-changing ReID, +11.7% mAP on COCAS+ real2 for clothes template based clothes-changing ReID when using only RGB images, +24.9% mAP on COCAS+ real2 for our newly defined language-instructed ReID, +4.3% on LLCM for visible-infrared ReID, +2.6% on CUHK-PEDES for text-to-image ReID. The datasets, the model, and code are available at https://github.com/hwz-zju/Instruct-ReID.
@@ 1 Introduction
Identifying individuals exhibiting significant appearance variations from multi-model descriptions is a fundamental aspect of human intelligence with broad applications [38, 35, 37, 73]. To imbue our machine with this capability, person re-identification (ReID) [78, 76, 8] has been introduced to retrieve images of the target person from a vast repository of surveillance videos or images across locations and time [11, 53]. Recently, significant advancements have been made in developing precise and efficient ReID algorithms and establishing benchmarks covering various scenarios, such as traditional ReID [75, 6, 54, 77], clothes-changing ReID (CC-ReID) [21, 13, 24, 46, 18], clothes template based clothes-changing ReID (CTCC-ReID) [66, 32], visible-infrared ReID (VI-ReID) [36, 57, 58, 68] and text-to-image ReID (T2I-ReID) [3, 4, 79, 30, 65]. However, focusing solely on one specific scenario possesses inherent limitations. For instance, customers must deploy distinct models to retrieve persons according to the query information, which significantly increases the cost for model training and results in inconvenience for the application. To facilitate real-world deployment, there is a pressing need to devise one generic framework capable of re-identifying individuals across all scenarios mentioned above.
@@
This paper proposes a new multi-purpose instruct-ReID task where existing 6 ReID settings can be formulated as its special cases. Specifically, the instruct-ReID uses query images and multi-model instructions as the model inputs, requiring the model to retrieve the same identity images from the gallery following the instructions. Using language or image instructions, models trained on the instruct-ReID task can be specialized to tackle diverse ReID tasks (Fig. 1a). For example, the clothes-changing ReID can be viewed as using the instruction “Ignore clothes” to retrieve. As another example, clothes template-based clothes-changing ReID can utilize a clothes template image as instruction. The proposed instruct-ReID task offers three significant advantages: easy deployment, improved performance, and easy extension to new ReID tasks. First, it enables cost-effective and convenient deployment in real-world applications. Unlike existing ReID approaches limited to specific tasks, instruct-ReID allows for utilizing a single model for various ReID scenarios, which is more practical in real applications. Second, as all existing ReID tasks can be considered special cases of instruct-ReID, we can unify the training sets of these tasks to exploit the benefits of more data and diverse annotations across various tasks – leading to enhanced performance. Third, instruct-ReID addresses a new language-instructed ReID task which requires the model to retrieve persons following language instructions. This setting is very practical in real applications for it enables customers to retrieve images they are particularly interested in through language descriptions. For example, customers can retrieve a woman wearing a black coat using one of her pictures and language instructions “Wear black coat”.
@@
To facilitate research in instruct-ReID, we introduce a new benchmark called OmniReID derived from 12 datasets  [^1] The discussion on ethical risks is provided in supplementary materials. of 6 different ReID tasks. The OmniReID benchmark exhibits two appealing characteristics. First, it emphasizes diversity by incorporating images from various domains, including surveillance scenarios and synthetic games. The diversity ensures that the trained models are robust and can effectively handle ReID tasks in various real-world scenarios. Second, the OmniReID benchmark achieves comprehensiveness by offering evaluation datasets to assess various ReID tasks comprehensively, which facilitates evaluating the generalization ability of diverse ReID methods.
@@
Based on OmniReID, we design a method with the proposed adaptive triplet loss for instruct-ReID. The typical triplet loss [44] only defines positive/negative pairs by identities, failing to align with instructions. Therefore, we propose a novel adaptive triplet loss to learn a metric space that preserves identity and instruction similarities. Specifically, we design an adaptive margin between two query-instruction pairs based on instruction similarities to pull features with similar instructions close and push features with different instructions apart. This loss incorporates the instruction information into the features representation and optimize the model to learn a metric space where similar instructions lead to closer features of query-instance pairs.
@@
In summary, the contributions of this paper are three folds. (1) We propose a new instruct-ReID task, where existing traditional ReID, clothes-changing ReID, clothes template based clothes-changing ReID, visible-infrared ReID, text-to-image ReID andlanguage-instructed ReID can be viewed as special cases. (2) To facilitate research on instruct-ReID, we establish a large-scale and comprehensive OmniReID benchmark consisting 12 publicly available datasets. (3) We propose an adaptive triplet loss to supervise the feature distance of two query-instruction pairs to consider identity and instruction alignments. Our method consistently improves previous models on 10 datasets of 6 ReID tasks. For example, our method improves +7.7%, +0.6%, +0.5% mAP on CUHK03, MSMT17, Market1501 for traditional ReID, +6.4%, +11.2%, +7.1% mAP on PRCC, LTCC, VC-Clothes for clothes-changing ReID when using RGB images only, +11.7% mAP on COCAS+ real2 for clothes template based clothes changing ReID, +4.3% mAP on LLCM for visible-infrared ReID, +2.6% mAP on CUHK-PEDES for text-to-image ReID, +24.9% mAP on COCAS+ real2 for the new language-instructed ReID.
@@ 2 Related work
Person Re-identification. Person re-identification aims to retrieve the same images of the same identity with the given query from the gallery set. To support the ReID task on all-weather application, various tasks are conducted on the scenarios with changing environments, perspectives, and poses [64, 6, 76, 4, 79, 80, 48]. Traditional ReID mainly focuses on dealing with indoor/outdoor problems when the target person wears the same clothes. To extend the application scenarios, clothes-changing ReID (CC-ReID) [66] and clothes template based clothes-changing ReID (CTCC-ReID) [32] are proposed. While CC-ReID forces the model to learn clothes-invariant features [21, 13], CCTC-ReID further extracts clothes template features [32] to retrieve the image of the person wearing template clothes. To capture person’s information under low-light environments, visible-infrared person ReID (VI-ReID) methods [68, 57, 58, 68] retrieve the visible (infrared) images according to the corresponding infrared (visible) images. In the absence of the query image, GNA-RNN [30] introduced the text-to-image ReID (T2I-ReID) task, which aims at retrieving the person from the textual description. However, existing researches focus on a single scenario, making it difficult to address the demands of cross-scenario tasks. In this paper, we introduce a new Instruct-ReID task, which can be viewed as a superset of the existing ReID tasks by incorporating instruction information into identification.
@@
Instruction Tuning. Instruction Tuning was first proposed to enable language models to execute specific tasks by following natural language instructions. Instruction-tuned models, e.g., FLAN-T5 [7], InstuctGPT [41]/ChatGPT [40], UPT [17] can effectively prompt the ability on zero- and few-shot transfer tasks. A few works borrowed the idea from language to vision. Flamingo [2], BLIP-2 [27], and KOSMOS-1 [20] learning with image-text pairs also show promising generalization on visual understanding tasks. While these methods aim to generate convincing language responses following the image or language instructions, we focus on retrieving the correct person following the given instructions by tuning a vision transformer.
@@
Multi-model Retrieval. Multi-model retrieval is widely used to align information from multiple modalities and improve the performance of applications. In multi-model retrieval, unimodal encoders always encode different modalities for retrieval tasks. For instance, CLIP [43], VideoCLIP [56], COOT [12] and MMV [1] utilize contrastive learning for pre-training. Other techniques like HERO [28], Clipbert [25], Vlm [55], TAM [39] and MVLT [22] focus on merging different modalities for retrieval tasks to learn a generic representation. Although there have been numerous studies on multi-modal retrieval, most are concentrated on language-vision pretraining or video retrieval, leaving the potential of multi-modal retrieval for person ReID largely unexplored. This paper aims to investigate this underexplored area to retrieve anyone with information extracted from multiple modalities.
@@ 3 OmniReID Benchmark
Language Annotation Generation. To generate language instruction for language-instructed ReID, we annotate COCAS+ Real1 and COCAS+ Real2 with language description labels. Similar to Text-to-Image ReID datasets, language annotations in OmniReID are several sentences that describe the visual appearance of pedestrians. We divide our annotation process into pedestrian attribute generation and attribution-to-language transformation.
@@ 4 IRM: Instruct ReID Model
As shown in Fig. 3 (a), the proposed Instruct ReID model consists of four parts: instruction generation (Sec. 4.1) for various ReID tasks, an editing transformer $\mathcal{E}_{e}$ (Sec. 4.2) and an instruction encoder $\mathcal{E}_{i}$ which is a visual language model, and a cross-model attention module $\mathcal{E}_{f}$. Given an instruction $\mathbf{T}$ associated with a query image $\mathbf{I}$, we obtain instruction features $\mathbf{F}_{\mathbf{T}}$ using the instruction encoder $\mathcal{E}i$. These extracted instruction features $\mathbf{F}_{T}$, along with query image tokens, are fed into the designed editing transformer $\mathcal{E}_{e}$ to obtain features $\mathbf{F}$ edited by instructions. Furthermore, we introduce an attention module $\mathcal{E}_{f}$ to efficiently combine features of the query image and instruction through the attention mechanism in the stacked transformer layers. Finally, the overall loss function is introduced (Sec. 4.4) for all ReID tasks, which includes our newly proposed adaptive triplet loss $\mathcal{L}_{atri}$ (Sec. 4.3) instead of traditional triplet loss.
@@ 4.1 Instruction Generation
Traditional ReID: Following Instruct-BLIP [10], we generate 20 instructions [^3] See all available instruction sentences in supplementary materials. from GPT-4 and randomly select one, e.g., “Do not change clothes.”, when training. The model is expected to retrieve images of the same person without altering image attributes, such as clothing. {mdframed}[backgroundcolor=gray!10] ### Query image:{Query image}
### Instruction: “Do not change clothes.”3
### Target image:{Output} Clothes-changing ReID: Simiar to Traditional ReID, the instruction is the sentence chosen from a collection of 20 GPT-4 generated sentences3, e.g., “With clothes changed”. The model should retrieve images of the same person even when wearing different outfits. {mdframed}[backgroundcolor=gray!10] ### Query image:{Query image}
### Instruction: “With clothes changed.”3
### Target image:{Output} Clothes template based clothes-changing ReID: The instruction is a clothes template for a query image while a cropped clothes image for a target image. The model should retrieve images of the same person wearing the provided clothes. We provide more examples for training and test in the supplementary materials. {mdframed}[backgroundcolor=gray!10] ### Query image:{Query image}
### Instruction:{Any clothes template image}
### Target image:{Output} Visual-Infrared ReID: The instruction is the sentence chosen from a collection of 20 GPT-4 generated sentences3, e.g., “Retrieve cross modality image”. The model should retrieve visible (infrared) images of the same person according to the corresponding infrared (visible) images. {mdframed}[backgroundcolor=gray!10] ### Query image:{Query image}
### Instruction: “Retrieve cross modality image.”3
### Target image:{Output} Text-to-Image ReID: The instruction is the describing sentences, and both images and text are fed into IRM during the training process. While in the inference stage, the image features and instruction features are extracted separately.  [^4] Image and instruction features are extracted separately in test stage. {mdframed}[backgroundcolor=gray!10] ### Image:{Image}4
### Instruction:{Sentences describing pedestrians}4
@@
Language-instructed ReID: The instruction is several sentences describing pedestrian attributes. We randomly select the description languages from the person images in gallery and provide to query images as instruction. The model is required to retrieve images of the same person described in the provided sentences. We provide more examples for training and test in the supplementary materials. {mdframed}[backgroundcolor=gray!10] ### Query image:{Query image}
@@ 4.2 Editing Transformer
where $\mathbf{F}_{\mathbf{T}}=\mathcal{E}_{i}(\mathbf{I})$ is the instruction feature extracted by $\mathcal{E}_{i}$ and $\mathbf{F}_{l-1}$ is the output feature of $(l\!-\!1)$-th zero-init transformer layer. The initial input $\mathcal{F}_{0}$ of the first layer is defined as $\mathbf{F}_{0}=[\mathbf{f}_{0}^{\text{CLS}},\mathbf{f}_{0}^{1},\mathbf{f}_{0}^{2},...,\mathbf{f}_{0}^{N}],$ where $\mathbf{f}^{\text{CLS}}_{0}$ is the [CLS] token, $(\mathbf{f}_{0}^{1},\mathbf{f}_{0}^{2},...,\mathbf{f}_{0}^{N})$ are the patch tokens of the query image and $N$ is the number of patches of the query image.